# Towards Single-Lens Controllable Depth-of-Field Imaging via All-in-Focus Aberration Correction and Monocular Depth Estimation

### v1

@@ IV-C Depth-aware PSF Simulation
TABLE I: Quantitative comparison of image restoration results between the two training schemes.
@@
IV-D Results of Synthetic Images Restoration
@@
TABLE II: Quantitative results of our depth-aware mechanism under the same DA2T scheme.
@@
As shown in Table I, the numerical results show that compared with the baseline, the model trained by the DA2T scheme has a stable improvement on four different models on two MOS. This proves that the scheme does not rely on a certain model and generalizes well to various architectures with consistent gains. On the MOS-S1 simulation data, the model trained using the DAT scheme outperforms the baseline by a margin of $0.398dB{\sim}0.6518dB$ on the NYUv2 dataset, and by $0.0797dB{\sim}0.5452dB$ on the Middlebury2014 dataset. On the MOS-S2 simulation dataset, the DAT-trained model demonstrates superior performance, surpassing the baseline by $3.0699dB{\sim}3.5300dB$ and $1.4089dB{\sim}2.0814dB$. It can be observed that, compared with MOS-S1, the performance gap between the models trained by the two training strategies is more significant in MOS-S2. This is because the PSFs in the MOS-S2 dataset change more dramatically with the scene depth. These experimental results indicate that the model trained using the DA2T scheme effectively restores degraded scenes with varying depths.
@@
Figure 9: Qualitative results on the real-world dataset. We constructed a scene with diverse objects positioned at varying depths within an optical laboratory under constant lighting conditions to capture a real-world dataset. The real-world dataset is captured using MOS-S1, which has also been employed to evaluate pre-trained CAC models. The DA2T scheme and depth-aware mechanism effectively restore objects at different depths.
@@
Restoration experiment of depth-aware mechanism. Similar to the experiment in the previous section, we sequentially replaced the RACM in DACN with four representative state-of-the-art super-resolution models, to investigate the recovery capacity of our proposed depth-aware mechanisms (RDICAB and D2CB). The quantitative results are shown in Table II. Across the four different network architectures, the designed depth-aware mechanisms exhibit stable improvements in the evaluation metrics. On the MOS-S1 dataset, the two proposed plug-and-play blocks, RDICAB and D2CB, lead to PSNR improvements ranging from $0.1780dB$ to $0.3317dB$. On the MOS-S2 dataset, the PSNR improvements brought by these two depth-aware mechanism blocks range from $0.1696dB$ to $0.8809dB$.
@@
As shown in Fig. 8, the method equipped with the depth-aware mechanism contributes to a superior visual reconstruction effect, faithfully recreating details closer to the ground truth and producing aberration-free images. Overall, our proposed two plug-and-play blocks (RDICAB and D2CB) demonstrate consistent and pronounced improvements in both quantitative metrics and visual quality across various architectures, showcasing the powerful performance and robust generalization capabilities of the depth-aware mechanism for aberration correction.
IV-E Results of Real-world Images Restoration
@@
The qualitative results of a real-world scene are displayed in Fig. 9, where we set up a scene in an optical laboratory under constant lighting conditions, with objects placed at varying depths. When restoring real-world data, we utilize RSTB as the RACM component in our proposed DACN, as it demonstrates the best performance on the simulated dataset. The DA2T model equipped with the depth-aware mechanism demonstrates superior visual results compared to the baseline (non-DA2T model without the depth-aware mechanism). By zooming in on image patches at different depths, it can be observed that our method successfully restores the texture information and reduces artifacts.
IV-F Results of Controllable Depth-of-field Image
@@
Figure 11: PSF estimation results. We train the Omni-Lens-Field network to represent the 4D PSFLib of four different lenses. The example demonstrates the PSF predicted by our network and ray-tracing PSF at three depths within the same FoV for the four lenses. The PSF produced by our method closely matches the GT.
@@
Figure 12: Visual results of controllable DoF imaging. (a) Aberration image simulated by MOS-S1. (b) Predicted depth map. (c) Predicted AiF image restored by our DACN. (d) Controllable DoF imaging of different lenses. We select the depth threshold to keep the “flower” clear and create the bokeh effect of three lenses, distinct from MOS-S1, beyond the threshold range.
@@
In our experiment, we aim to use the Omni-Lens-Field model to characterize the 4D PSFLib of four different lenses: the MOS-S1, MOS-S2, a Double Gauss lens, and a 6P lens automatically designed by OptiFusion [59]. In Fig. 10, we explore the selection of the number of hidden layers $N$, to find the balance between storage space and performance. Based on these results, we set $N{=}5$, which allows for precise PSF prediction within a relatively small storage size. Compared to the original 4D PSFLib of four lenses, which requires $67.8GB$ of storage space, our Omni-Lens-Field model achieves the same representational capacity while occupying only $0.24GB$ of storage. This represents a $282.5$-fold reduction in the storage space requirements. Additionally, the performance of our method, as measured by PSNR and SSIM on the validation set, reaches $49.2855dB$ and $0.9953$ respectively, which confirms that our method is storage-efficient and achieves high-quality prediction with little storage consumption. In Fig.11, we present the estimated PSF results of four lenses at depths of $1m$, $2m$, and $5m$ with a FoV of $0.5$. At each depth, our predictions closely match the ray-tracing PSFs. The detailed optical parameters of these four lenses and more prediction results are shown in the Appendix.
@@
For a degraded image captured by the MOS, our AiF aberration correction method can recover an AiF aberration-free image that achieves wide DoF imaging. Based on the restored AiF image, by leveraging the predicted depth map and the depth-aware PSF map inferred by Omni-Lens-Field, an image can be obtained where the main subject remains distinctly visible, while the background is given the bokeh effect of any desired lens, thereby achieving a shallow DoF effect of different lenses, as depicted in Eq. (12). Note that both the subject and the range to be kept sharp are user-controlled. In this way, controllable DoF imaging is realized. The visual results are displayed in Fig. 12.
IV-G Ablation Study
To verify the effectiveness of our two plug-and-play modules, RDICAB and D2CB, we use RB [23] as our RACM module and conduct ablation experiments on NYUv2 and Middlebury2014 datasets. All models are trained in the same experimental environment and settings. As proofed in Table III, both RDICAB and D2CB are beneficial to improving the aberration image recovery ability of the model, with RDICAB achieving an increase of $0.0864dB$ in PSNR, while D2CB resulting in an increase of $0.1328dB$.
TABLE III: Ablation study on RDICAB and D2CB.
@@ V Conclusion
Limitations and future work directions. In aberration correction, the focus distance has not been fully considered; instead, the focus distance is fixed. This is because, even with adjustable focusing cameras, the focus plane is only positioned at a selected depth during capture, rather than across all depths. However, it is essential to consider the focus distance to achieve the imaging effect of a true high-end lens with a single lens, which requires simulating a 5D PSFLib. Unfortunately, this significantly increases the computational cost and storage requirements of ray tracing. In the future, we intend to utilize a 5D PSFLib for controllable DoF imaging with a single lens which remains a topic worth exploring.
@@ Appendix A Lens Data
TABLE A.2: Lens data for the MOS-S1 used in this paper.
@@
TABLE A.3: Lens data for the Double Gauss Lens used in this paper.
@@
TABLE A.4: Lens data for the 6PLENS used in this paper. This lens is automatically designed by OptiFusion [59]
@@
Appendix B Aberration Image Simulation
In our experiments, we need degraded-sharp image pairs and the corresponding GT depth map. Unfortunately, the absence of such datasets prompts us to simulate our own required training data. We select the NYUv2 [67] and Middlebury2014 [68] datasets, which provide RGBD images. The resolution ratio of training and testing images is dictated by the camera sensor resolution. Although optical aberrations are more pronounced in high-resolution images, processing ultra-high-resolution images produced by megapixel commercial cameras is impractical during network training. Fortunately, due to the physical properties of the PSF, downsampling is a viable solution. Therefore, we strike a balance between performance and memory consumption. The sensor size is $4000{\times}6000$, we downsample it to $1280{\times}1920$. The image resolution of the NYUv2 dataset is $480{\times}640$, so we consider these images as part of images with a resolution of $1280{\times}1920$. Specifically, we randomly generate an index for each image and select the corresponding depth-aware PSF map according to the index for patch-wise convolution to generate aberrated images. The simulation process is shown in Fig. B.1. The image resolution in the Middlebury2014 dataset is reshaped and cropped to $1280{\times}1920$.
Figure B.1: Depth-aware aberration image dataset simulation process.
Appendix C More PSF Estimation Results
@@
In this section, we show the PSFs of four lenses estimated by the Omni-Lens-Field model at different fields of view and depths, including more results for two fields of view ($0.0$ and $1.0$) and three depths ($1m$, $2m$, and $5m$). As illustrated in Fig. C.2, our Omni-Lens-Field model can predict PSFs at different fields of view and depths close to GT with high quality, which shows that the 4D PSF library of various lenses can be effectively represented by our model.
@@
Figure C.1: 2D views of the four used lenses. (a) The MOS-S1 has a focal length of $20.0mm$, an F-number of $5.0$, and a field of view of $42.6^{\circ}$. (b) The MOS-S2 has a focal length of $24.7mm$, an F-number of $3.5$, and a field of view of $34.0^{\circ}$. (c) The Double Gauss Lens has a focal length of $4.0mm$, an F-number of $1.0$, and a field of view of $2.5^{\circ}$. (d) The 6P Lens has a focal length of $26.44mm$, an F-number of $1.8$, and a field of view of $32.0^{\circ}$.
@@
Figure C.2: More PSF estimation results. This figure shows the PSF prediction results of our model at two FoVs (00 and $1$) and three depths ($1m$, $2m$, $5m$). The PSF predicted by our model is close to GT.

### v2

@@ IV-C Depth-aware PSF Simulation
TABLE II: Quantitative comparison of image restoration results between the two training schemes.
@@
IV-D Results of Monocular Depth Estimation
We present the depth estimation results, where the GT images and aberration images are predicted using the same metric MDE model [11].
As shown in Table III, quantitative experimental results indicate that the depth predicted from clear images is not always superior to that predicted from aberration images. On the Middlebury2014 dataset [70], the depth predicted from aberration images can even outperform that from clear images. Therefore, aberration does not consistently have a negative impact on depth prediction; in some cases, it can implicitly encode depth information, thereby enhancing the accuracy of the predictions [18, 19, 20, 21].
It is worth noting that the distortion introduced by the optical system does not impact the overall image quality. Since this paper focuses on aberration correction, we ensure pixel alignment between the GT image and the aberration image by applying the same optical system distortion to both during simulation. Consequently, the GT images vary between different optical systems, which explains the discrepancies in the quantitative metrics of the depth maps predicted by the same MDE model for GT images from two different optical systems. Compared to the depth predictions from GT images, the depth predictions from aberration images still provide reliable depth information. The qualitative results are shown in the appendix.
TABLE III: Quantitative comparison of monocular depth estimation using the same model on GT images and aberration images.
Lens name Method Image Dataset
NYUv2 MiddleBury2014
AbsRel $\downarrow$ RMSE $\downarrow$ $\delta>1.25\uparrow$ AbsRel $\downarrow$ RMSE $\downarrow$ $\delta>1.25\uparrow$
MOS-S1 Unidepth [11] GT 0.082 0.328 0.955 0.429 1.143 0.321
Aberration 0.096 0.366 0.934 0.425 1.179 0.273
MOS-S2 GT 0.076 0.311 0.961 0.477 1.292 0.229
Aberration 0.097 0.369 0.930 0.442 1.254 0.300
IV-E Results of Synthetic Images Restoration
@@
TABLE IV: Quantitative image restoration results of our depth-aware mechanism under the same DA2T scheme.
@@
Figure 9: Qualitative results on the real-world indoor dataset. We constructed a scene with diverse objects positioned at varying depths within an optical laboratory under constant lighting conditions to capture a real-world indoor dataset. The real-world indoor dataset is captured using MOS-S1, which has also been employed to evaluate pre-trained CAC models. The DA2T scheme and depth-aware mechanism effectively restore objects at different depths.
@@
As shown in Table II, the numerical results show that compared with the baseline, the model trained by the DA2T scheme has a stable improvement on four different models on two MOS. This proves that the scheme does not rely on a certain model and generalizes well to various architectures with consistent gains. On the MOS-S1 simulation data, the model trained using the DAT scheme outperforms the baseline by a margin of $0.398dB{\sim}0.6518dB$ on the NYUv2 dataset, and by $0.0797dB{\sim}0.5452dB$ on the Middlebury2014 dataset. On the MOS-S2 simulation dataset, the DA2T-trained model demonstrates superior performance, surpassing the baseline by $3.0699dB{\sim}3.5300dB$ and $1.4089dB{\sim}2.0814dB$. It can be observed that, compared with MOS-S1, the performance gap between the models trained by the two training strategies is more significant in MOS-S2. This is because the PSFs in the MOS-S2 dataset change more dramatically with the scene depth. These experimental results indicate that the model trained using the DA2T scheme effectively restores degraded scenes with varying depths.
@@
Figure 10: Qualitative results on the real-world outdoor dataset. The real-world outdoor dataset is captured using MOS-S1, which has also been employed to evaluate pre-trained CAC models. The proposed training scheme and depth-aware mechanism effectively restore outdoor scenes with dynamic depth variations.
@@
Restoration experiment of depth-aware mechanism. Similar to the experiment in the previous section, we sequentially replaced the RACM in DACN with four representative state-of-the-art super-resolution models, to investigate the recovery capacity of our proposed depth-aware mechanisms (RDICAB and D2CB). The quantitative results are shown in Table IV. Across the four different network architectures, the designed depth-aware mechanisms exhibit stable improvements in the evaluation metrics. On the MOS-S1 dataset, the two proposed plug-and-play blocks, RDICAB and D2CB, lead to PSNR improvements ranging from $0.1780dB$ to $0.3317dB$. On the MOS-S2 dataset, the PSNR gains brought by these two depth-aware mechanism blocks range from $0.1696dB$ to $0.8809dB$.
@@
As shown in Fig. 8, the method equipped with the depth-aware mechanism contributes to a superior visual reconstruction result, faithfully recreating details closer to the ground truth and producing aberration-free images. Overall, our proposed two plug-and-play blocks (RDICAB and D2CB) demonstrate consistent and pronounced improvements in both quantitative metrics and visual quality across various architectures, showcasing the powerful performance and robust generalization capabilities of the depth-aware mechanism for aberration correction.
IV-F Results of Real-world Images Restoration
@@
Indoor Scene Evaluation. The qualitative results of a real-world scene are displayed in Fig. 9, where we set up a scene in an optical laboratory under constant lighting conditions, with objects placed at varying depths. When restoring real-world data, we utilize RSTB as the RACM component in our proposed DACN, as it demonstrates the best performance on the simulated dataset. The DA2T model equipped with the depth-aware mechanism demonstrates superior visual results compared to the baseline (non-DA2T model without the depth-aware mechanism). By zooming in on image patches at different depths, it can be observed that our method successfully restores the texture information and reduces artifacts.
Outdoor Scene Evaluation. The qualitative results of outdoor scenes are displayed in Fig. 10. Real-world outdoor scenes, characterized by more dynamic depth variations, present greater challenges for restoration. We observed that the non-DAT model struggles to restore scenes with certain depth ranges, leading to unpleasant artifacts. In contrast, the DAT-trained model, equipped with a depth-aware mechanism, effectively reduces artifacts and delivers visually superior restoration results.
IV-G Results of Controllable Depth-of-Field Imaging
TABLE V: Inference time analysis of Depth Estimation Module, Computational Aberration Correction Modules, and Omni-Lens-Field. The inference time is calculated with the input resolution of $1280{\times}1920$ on a single 3090 GPU. The percentage represents the proportion of time that the Omni-Lens-Field occupies within the entire pipeline.
Method
Inference
Time (s)
Method
Inference
Time (s)
Method
Inference
Time (s)
Percentage
UniDepth [11] 0.212 RB [23] 0.092 Omni-Lens- Field 0.028 8.4%
RDG [26] 0.657 3.1%
PSA [25] 0.369 4.6%
RSTB [24] 0.752 2.8%
@@ IV-G Results of Controllable Depth-of-Field Imaging
Figure 12: PSF estimation results. We train the Omni-Lens-Field network to represent the 4D PSFLib of four different lenses. The example demonstrates the PSF predicted by our network and ray-tracing PSF at three depths within the same FoV for the four lenses. The PSF produced by our method closely matches the GT. The size of each PSF is $41{\times}41$ at a pixel size of $12.2~{}{\mu}m$.
@@
Figure 13: Qualitative results of controllable DoF imaging on the real-world outdoor dataset. (a) Aberration image simulated by MOS-S1. (b) Predicted depth map. (c) Predicted AiF image restored by our DACN. (d) Controllable DoF imaging of different lenses. We select the depth threshold to keep the “wall” clear and create the bokeh effects of three lenses, distinct from MOS-S1, beyond the threshold range.
@@
In our experiment, we aim to use the Omni-Lens-Field model to characterize the 4D PSFLib of four different lenses: the MOS-S1, MOS-S2, a Double Gauss lens, and a 6P lens automatically designed by OptiFusion [61]. In Fig. 11, we explore the selection of the number of hidden layers $N$, to find the balance between storage space and performance. Based on these results, we set $N{=}5$, which allows for precise PSF prediction within a relatively small storage size. Compared to the original 4D PSFLib of four lenses, which requires $67.8GB$ of storage space, our Omni-Lens-Field model achieves the same representational capacity while occupying only $0.24GB$ of storage. This represents a $282.5$-fold reduction in the storage space requirements. Additionally, the performance of our method, as measured by PSNR and SSIM on the validation set, reaches $49.2855dB$ and $0.9953$ respectively, which confirms that our method is storage-efficient and achieves high-quality prediction with little storage consumption. The time consumption of the model should also be taken into account. As shown in Table V, the Omni-Lens-Field accounts for at most 8.4% of the inference time in the entire pipeline, which is within an acceptable range.
In Fig.12, we present the estimated PSF results of four lenses at depths of $1m$, $2m$, and $5m$ with a FoV of $0.5$. At each depth, our predictions closely match the ray-tracing PSFs. The detailed optical parameters of these four lenses and more prediction results are shown in the Appendix.
@@
For a degraded image captured by the MOS, our AiF aberration correction method can recover an AiF aberration-free image that achieves wide DoF imaging. Based on the restored AiF image, by leveraging the predicted depth map and the depth-aware PSF map inferred by Omni-Lens-Field, an image can be obtained where the main subject remains distinctly visible, while the background is given the bokeh effect of any desired lens, thereby achieving a shallow DoF effect of different lenses, as depicted in Eq. (12). Note that both the subject and the range to be kept sharp are user-controlled. In this way, controllable DoF imaging is realized. The real-world outdoor controllable DoF image results are presented in Fig. 13.
IV-H Ablation Study
To verify the effectiveness of our two plug-and-play modules, RDICAB and D2CB, we use RB [23] as our RACM module and conduct ablation experiments on NYUv2 and Middlebury2014 datasets. All models are trained in the same experimental environment and settings. As proofed in Table VI, both RDICAB and D2CB are beneficial to improving the aberration image recovery ability of the model, with RDICAB achieving an increase of $0.0864dB$ in PSNR, while D2CB resulting in an increase of $0.1328dB$.
TABLE VI: Ablation study on RDICAB and D2CB.
@@ IV-H Ablation Study
Additionally, we further investigate the assignments of Query ($Q$), Key ($K$), and Value ($V$) in RDICAB. In Table. VII, the CAC model, which obtains queries from the image branch and keys and values from the depth branch, achieves better aberration image restoration results. Under both assignment conditions, we ensure that the long skip residual connection in RDICAB uses image features. As shown in Eq. 8, assigning $Q$ to the image branch while assigning $K$ and $V$ to the depth branch enables the cross-attention mechanism to generate depth information modulated by image features. This modulated depth information is then added to the image features and fed into subsequent processing modules. The added depth information significantly enhances the model’s ability to handle depth-varying degradations. In contrast, reversing this assignment results in the generation of single-modal image features, which fail to provide additional information to the subsequent modules, thereby limiting the overall performance of the model.
Finally, to evaluate the impact of depth maps predicted by aberration and GT images on the CAC task, we conduct comparative experiments using each as the depth branch input for the CAC model. As illustrated in Table VIII, the two different depth branch inputs have minimal impact on the model’s aberration correction performance. This suggests that directly using a pre-trained MDE model to predict depth from aberration images can be effectively applied to downstream tasks, as the predicted depth maps are reliable.
TABLE VII: Ablation study on the assignments of Query ($Q$), Key ($K$), and Value ($V$).
Method $Q$ $K$ $V$ NYUv2 MiddleBury2014
[…]
RB [23] Image Depth Depth 34.5264 0.9428 0.0723 35.2610 0.9443 0.1052
Depth Image Image 34.4449 0.9422 0.0728 35.1272 0.9430 0.1077
TABLE VIII: Ablation study on the impact of depth predicted by GT images and aberration images on the CAC task.
Method Image NYUv2 MiddleBury2014
[…]
RB [23] GT 34.6373 0.9437 0.0701 35.4822 0.9447 0.1041
Aberration 34.6096 0.9440 0.0698 35.5162 0.9445 0.1021
@@ V Conclusion
Limitations and future work directions. At the dataset level, the depth range of the scene that our method can effectively restore is limited to a certain extent by the depth range of the training dataset. In future research, creating a paired RGB and depth dataset with a wide variety of scene content and a broader depth range will be crucial for enhancing the applicability of our method. In ray tracing simulation, the influences of the scattering media during the light transmission process are not considered. This paper proposes the first framework for single-lens controllable DoF imaging, focusing on common application scenarios, where the scattering has little impact on the simulation results. However, when generalizing this framework to other extreme scenarios (such as foggy weather), it is necessary to consider the scattering characteristics of the medium. In aberration correction, the focus distance has not been fully considered; instead, the focus distance is fixed. This is because, even with adjustable focusing cameras, the focus plane is only positioned at a selected depth during capture, rather than across all depths. However, it is essential to consider the focus distance to achieve the imaging effect of a true high-end lens with a single lens, which requires simulating a 5D PSFLib. This, unfortunately, significantly increases the computational overhead and incurs cost-prohibitive storage requirements of ray tracing. In the future, we intend to utilize a 5D PSFLib for controllable DoF imaging with a single lens.
@@ Appendix A Lens Data
TABLE A.2: Lens data for the MOS-S1 used in this work.
@@
TABLE A.3: Lens data for the Double Gauss Lens used in this work.
@@
TABLE A.4: Lens data for the 6PLENS used in this work. This lens is automatically designed by OptiFusion [61]
@@
Appendix B Impact of Diffraction Effects on Ray Tracing Simulation
We explain the reasons why diffraction effects can be ignored in our prototype camera (MOS-S1) from the following three perspectives:
First, the aperture stop of the MOS-S1 has a diameter of approximately $4mm$, which greatly exceeds the wavelength range of visible light, from $380nm$ to $700nm$. As a result, since the diameter of the aperture stop is much larger than the wavelength of light, no significant diffraction occurs at the aperture stop.
Besides, MOS-S1 utilizes a spherical lens with a continuously varying surface spherical height. There are no microstructures akin to diffraction elements (such as metasurfaces or diffraction optical elements, DOE), which means that no diffraction effects occur on the optical surface.
Lastly, following your suggestions, we also calculate the Airy disk of this optical system and compare it with the pixel size. The Airy disk calculation formula is as follows:
$d_{Airy}=2.44\cdot\lambda\cdot\text{F\#},$# , (B.1)
where F# denotes image-space F-number, and $\lambda$ denotes visible light wavelength. Taking ${\lambda}{=}550nm$ in the formula, the diameter of the Airy disk is approximately $6.7{\mu}m$. When ray tracing the PSF simulation, the image resolution is set to $1280{\times}1920$, and the physical diagonal size of the CMOS sensor is $28.2mm$, so the physical size of each pixel is about $12.2{\mu}m$. Therefore, the Airy disk diameter is smaller than the physical size of a pixel and the diffraction effect can be ignored.
In summary, it is reasonable to ignore the influence of diffraction effects in our prototype camera.
Appendix C Depth-aware Degradation-adaptive Training pipeline
In order to embed the depth-varying optical degradation into the network training, the DA2T scheme integrates the depth-aware image simulation module into the training process, as illustrated in Fig. C.1(c). GT image and its corresponding depth map are used as input during the overall pipeline (Fig. C.1(a)).
According to the depth map, the corresponding PSF is retrieved from the PSF library for each patch, and a depth-aware PSF map is formed based on the depth value, which is used to generate the aberration image. The entire training process enables the network to adaptively learn to handle the optical degradation process that varies with depth.
Figure C.1: Depth-aware Degradation-adaptive Training pipeline. (a) GT image and corresponding depth map are given as the input. (b) based on the depth map, a corresponding PSF is searched in the PSF library for each patch according to the depth value, forming a depth-aware PSF map (purple path). (c) the GT image and the corresponding PSF map are processed using patch-wise convolution to generate the aberration image. Distortion simulation is then both the GT image and the aberration image, which are used as the training data. (pink path). (d) the CAC model takes the aberration image to estimate the recovered image (green path).
Appendix D More PSF Estimation Results
@@
In this section, we show the PSFs of four lenses estimated by the Omni-Lens-Field model at different fields of view and depths, including more results for two fields of view ($0.0$ and $1.0$) and three depths ($1m$, $2m$, and $5m$). As illustrated in Fig. D.2, our Omni-Lens-Field model can predict PSFs at different fields of view and depths close to GT with high quality, which shows that the 4D PSF library of various lenses can be effectively represented by our model.
@@
Figure D.1: 2D views of the four used lenses. (a) The MOS-S1 has a focal length of $20.0mm$, an F-number of $5.0$, and a field of view of $42.6^{\circ}$. (b) The MOS-S2 has a focal length of $24.7mm$, an F-number of $3.5$, and a field of view of $34.0^{\circ}$. (c) The Double Gauss Lens has a focal length of $4.0mm$, an F-number of $1.0$, and a field of view of $2.5^{\circ}$. (d) The 6P Lens has a focal length of $26.44mm$, an F-number of $1.8$, and a field of view of $32.0^{\circ}$.
@@
Figure D.2: More PSF estimation results. This figure shows the PSF prediction results of our model at two FoVs (00 and $1$) and three depths ($1m$, $2m$, $5m$). The PSF predicted by our model is close to GT.
Appendix E Quantitative Analysis of Different state-of-the-art MDE models on the DAMOS Dataset
In this section, we compare the depth prediction performance of four state-of-the-art metric MDE models on the DAMOS dataset: ZoeDepth [76], Metric3Dv2 [77], DepthAnythingV2 [78], and Unidepth [11].
To verify the robustness of these metric depth estimation methods, we evaluate on the NYUv2 [69] and Middlebury2014 [70] datasets of two lenses. Following previous studies [76, 77, 11], we use metrics such as absolute relative error (AbsRel), accuracy under threshold ($\delta_{i}<1.25^{i},\,i=1,2,3$), root mean squared error (RMSE), root mean squared error in log space ($\text{RMSE}_{\log}$).
Table E.1 clearly shows Unidepth consistently ranks in the top two across all datasets, demonstrating strong zero-shot generalization capability. All of the above models use their best-performing versions as presented in the paper.
Based on the above quantitative results, we choose UniDepth as the model to obtain the metric depth information of the scene.
TABLE E.1: Quantitative comparison of monocular depth estimation on aberration images using different models. “*” means that the corresponding camera intrinsics are required for inference. “†” means using its metric depth estimation model.
Lens name Method NYUv2 Middlebury2014
[…]
MOS-S1 ZoeDepth [76] 0.886 0.974 0.991 0.113 0.443 0.156 0.377 0.609 0.746 0.335 1.128 0.457
Metric3Dv2 [77] 0.759 0.946 0.978 0.156 0.624 0.229 0.145 0.375 0.764 0.828 2.106 0.578
DepthAnythingv2 [78] 0.791 0.960 0.984 0.167 0.621 0.201 0.130 0.283 0.612 0.465 1.345 0.678
UniDepth [11] 0.934 0.987 0.995 0.096 0.366 0.130 0.273 0.630 0.840 0.425 1.179 0.465
MOS-S2 ZoeDepth [76] 0.868 0.971 0.991 0.122 0.466 0.166 0.382 0.604 0.751 0.333 1.117 0.455
Metric3Dv2 [77] 0.788 0.949 0.976 0.144 0.581 0.223 0.089 0.313 0.700 0.928 2.381 0.625
DepthAnythingv2 [78] 0.712 0.948 0.983 0.204 0.671 0.219 0.122 0.306 0.636 0.464 1.338 0.682
UniDepth [11] 0.930 0.986 0.995 0.097 0.369 0.130 0.300 0.583 0.779 0.442 1.254 0.468